# Detecting Supernova Axions with IAXO

### v1

@@ 1 Introduction
Given the current technology capabilities, the detection of axions from a nearby SN event is a realistic possibility and the construction of a dedicated MeV detector for the next generation of axion helioscopes is highly motivated. First, a sensitivity to SN axions would allow us to explore regions of the axion parameter space not accessible through solar observations. Second, such a detection could shed light on important aspects of the SN physics, not accessible through other experimental tests. SNe are very rare events with current estimates predicting a rate of $1.63\pm 0.46$ per century [14]. It is, therefore, imperative to ensure the highest level of preparedness for such an eventuality.
@@ 2 Supernova axion production
If axions are coupled to nuclear matter, the main production channels in the hot and dense SN core depend on their coupling to nucleons $g_{aN}$, with $N=n,\,p$ for neutrons and protons respectively [18, 19, 20]. The first considered channel for axion emission from the SN nuclear medium is $NN$ bremsstrahlung $N+N\rightarrow N+N+a$ [21, 22, 23, 24, 25]. As depicted in the left panel of Fig. 1, in this process an axion is produced as a consequence of strong interactions between two nucleons in the medium. The state-of-the-art calculation for the bremsstrahlung emission rate has been performed in Ref. [26] and accounts for corrections beyond the usual one-pion exchange (OPE) approximation for the nuclear interaction potential. In particular, a non-vanishing mass for the exchanged pion [27], the contribution from the two-pion exchange [28], effective in-medium nucleon masses and multiple nucleon scattering effects [29, 30] have to be taken into account to obtain a reliable estimation of the bremsstrahlung emissivity. Finally, Ref. [31] has discussed the impact of finite-density effects over ALP-nucleon couplings in dense environments. Furthermore, the right panel of Fig. 1 displays a schematic representation of pionic Compton-like processes $\pi+N\rightarrow a+N$, in which a pion is converted into an axion after the scattering onto a non-relativistic nucleon. Since in the SN nuclear medium the fraction of thermal pions $Y_{\pi}$ was expected to be extremely small $Y_{\pi}\sim\mathcal{O}(10^{-4})$ [32], the contribution from pion conversions has been overlooked for a long time. Nevertheless, the authors of Ref. [33] pointed out that strong interactions can magnify the abundance of negatively-charged pions. Thus, the pion conversion emission rate has been re-estimated in Ref. [34], realizing that, under the assumption of Ref. [33], it could be comparable and even dominant over the bremsstrahlung contribution. Moreover, it has been recently shown that the contribution from the contact interaction term [35] and the $\Delta(1232)$ resonance [36] can significantly enhance axion emissivity via pion conversion. We address the reader to Refs. [18, 37] for the complete expressions of both $NN$ bremsstrahlung and pion conversion emission rates.
@@
If axions are weakly coupled to nuclear matter, ${g_{aN}\lesssim 10^{-8}}$, they are in the free-streaming regime, in which they can leave the star without being reabsorbed by the nuclear medium in the SN core. Therefore, they can act as an additional energy-loss channel during the SN cooling phase, contributing to the shortening of the neutrino burst duration [39]. In particular, consistency with observations of the neutrino burst from SN 1987A [^1] Current state-of-the-art SN simulations, including convection and updated neutrino-nucleon opacities, are in tension with late-time SN 1987A events, since they predict a cooling time shorter than the observed one [40]. requires ${g_{ap}\lesssim 5\times 10^{-10}}$ [19]. On the other hand, for sufficiently large couplings $g_{aN}\gtrsim 10^{-8}$, the SN core becomes “optically thick” for the produced axions, so that they can be also reabsorbed inside the core by means of the inverse processes $N+N+a\rightarrow N+N$ and $a+N\rightarrow\pi+N$ [19]. In this regime, dubbed trapping regime, axions are not able to escape the inner regions of the core and their emission can be approximately described as a blackbody radiation from a last-scattering surface [41], in close analogy with the case of neutrinos. We highlight that a complete treatment of the trapping regime would require the self-consistent addition of axions in SN simulations, since they could significantly impact the SN evolution. Relying just on the SN cooling argument, the range $5\times 10^{-10}\lesssim g_{ap}\lesssim 3\times 10^{-6}$ is excluded. However, larger values of the coupling can be constrained by the non observation of axion-induced events in Kamiokande-II during SN 1987A [42, 19]. Nevertheless, despite these constraints, even the trapped region remains a subject of interest for axion experiments [17].
@@
In the following, we adopt as benchmark SN model the 1D spherical-symmetric GARCHING group’s SN model SFHo-s18.8 provided in Ref. [38], used also in previous studies about SN axions (see, e.g. Refs. [43, 44, 45, 18, 19, 46, 20]). The simulation, based on the neutrino-hydrodynamics code PROMETHEUS-VERTEX [47], is launched from a stellar progenitor with mass $18.8~{}M_{\odot}$ [48] and leads to a neutron star (NS) with baryonic mass $1.35~{}M_{\odot}$, a typical value expected for a SN explosion event [15]. More details on the employed SN model are provided in Appendix A. The presence of pions in the SN core has been taken into account by following the procedure illustrated in Ref. [49], including the pion-nucleon interaction as described in Ref. [33]. Figure 2 shows the axion emission spectrum integrated over $10\,\operatorname{s}$ in both the free-streaming (black lines) and trapping regimes (red line) for our benchmark SN model. In the free-streaming regime, the spectrum is characterized by a bimodal shape, since the emission peak associated with the $NN$ bremsstrahlung is located at axion energies $E_{a}\sim 50\,\operatorname{MeV}$, while the pion conversion contribution peaks at $E_{a}\sim 200\,\operatorname{MeV}$ [18, 19]. This is related to the different nature of the two processes. In particular, the bremsstrahlung behaves as a quasi-thermal process, so that the average energy of the emitted ALP can be estimated as $E_{a}\simeq\frac{3}{2}T$  [46]. On the other hand, the pion conversion is highly non-thermal and the average ALP energy is given by $E_{a}\simeq\,m_{\pi}+\frac{3}{2}T$, where $m_{\pi}\simeq 135\,\operatorname{MeV}$ is the pion mass.
@@ 3 Direct detection of SN axions in IAXO
7 HD 157244 $\beta$ Arae $0.219\pm 0.010$ 17:25:17.99 -55:31:47.57
8 HD 39801 Betelgeuse/$\alpha$ Orionis 0.222 $\pm$ 0.040 05:55:10.31 +07:24:25.4
@@
It is evident that the detection of SN axions using IAXO and BabyIAXO relies heavily on our ability to predict core-collapse SN events with sufficient advance notice and angular precision to plan observations effectively. For this purpose, one has to rely on an early-alert system. Effectively, this is possible through the observation of pre-SN neutrinos, as the neutrino production rises rapidly during the last stages of nuclear burning [53]. Although the flux remains substantially lower—and at lower energies—than what is expected during the explosion itself, these neutrino fluxes are predicted to be observable in dedicated detectors minutes to days before the rise of the electromagnetic emission, providing the desired early alert (see, e.g., Ref. [53]). Dedicated analyses indicate, for example, that the combined analysis of KamLAND and Super-Kamiokande (SK) can detect the pre-SN neutrino flux from stars up to $\sim$ 500 pc, providing an early warning up to $\sim 12$ hours before the SN event [54]. Extracting directional information is more challenging. A detailed study of the potential to accurately detect the position of the source from pre-SN observations was presented in Ref. [50]. Information about the upcoming event, including positional data, is expected to be available hours before the event, a time more than sufficient to turn the IAXO telescope towards the star. The directional sensitivity of these events is relatively low, due to the weak forward-backward asymmetry of the inverse beta decay process used to detect the neutrino events. For current liquid scintillators, the low value of this asymmetry translates into a moderate angular resolution of around $60^{\circ}$ when 200 events are detected, which is realistic for a star like Betelgeuse. The use of a lithium-loaded liquid scintillator (LS-Li) would substantially increase the forward-backward asymmetry, pushing the angular resolution to about $15^{\circ}$ with the same number of events. Further improvements are possible with proposed future neutrino detectors, such as THEIA [55].
@@
Despite the moderate angular precision expected in the case of pre-SN neutrino events, the information provided by an early warning system may suffice to identify the SN candidate, given that the sample of possible close-by candidates is known (see Tab. 1), and that their relative angular distances are in most cases relatively large (cf. Tab. 5 in Appendix B).
@@ 4 Sensitivity prospects
$S_{\gamma}=g_{a\gamma}^{2}\,g_{aN}^{2}\int_{E_{i}}^{E_{f}}\mathrm{d}E\,\frac{\mathrm{d}n_{\gamma}}{\mathrm{d}E}\leq\operatorname{ln}(20)$ (4.5)
Sensitivity prospects are shown in Fig. 5 and 6 for two different SN progenitor candidates, namely Spica ($\alpha$ Virginis) and Betelgeuse ($\alpha$ Orionis). These prospects have been computed in different helioscope scenarios: BabyIAXO, consisting of two magnetic bores, with a length $L=10$ m and an area $A=0.77$ m2 as well as an average magnetic field $B=2$ T; IAXO, which will be equipped with 8 magnet bores with $L=20$ m and $A=2.3$ m2, providing a magnetic field of strength $B=2.5$ T; and IAXO+, which represents an enhanced version of IAXO, with a magnetic field $B=3.5$ T, an area $A=3.9$ m2 and magnet bores with $L=22$ m. These helioscope scenarios are described in Ref. [9] and summarized in Table 4. For the high-energy gamma detector we assume a conservative detection efficiency of $\epsilon=95\%$ and an energy threshold of 3 MeV.
@@
Figure 5: Sensitivity prospects for BabyIAXO, IAXO and IAXO+ as SNScopes for Betelgeuse (violet, left panel) and Spica (green, right panel) in the ALP free streaming regime considering bremsstrahlung emission and both bremsstrahlung and pion conversion emission. In the upper panels, we set the ALP-nucleon coupling to ${g_{aN}=10^{-9}}$, while in the lower panels we set it to the maximum value allowed by the the SN cooling argument ${g_{aN}=5\times 10^{-10}}$. The black transparent regions show the sensivity prospects of IAXO, BabyIAXO and IAXO+ as helioscopes.
@@
Instead of using the natural likelihood, it is more convenient to work with the logarithm of the likelihood $\ln{L}$. According to Wilks’ theorem [59], if certain general conditions are satisfied, the minimum of $-2\ln{L}$ approaches a $\chi^{2}$ distribution
@@
The likelihood analysis is performed for a fixed axion mass in the coherence region ($m_{a}=0.01$ eV) while scanning over the $g_{a\gamma}^{2}\times g_{aN}^{2}$ and $\delta$ parameter space. The minimum of the $\chi^{2}$ distribution is extracted, in which the shift of one $\chi^{2}$ unit corresponds to one standard deviation ($\sigma$) in the parameter estimation. Fig. 7 and Fig. 8 show an illustrative example of the likelihood analysis results for 10 detected counts, using as input distribution the free streaming case of Fig. 2 with $\delta=1$ and a coupling constant of $g_{aN}=5\times 10^{-10}$. Being the best fit parameters close to their truth values, these results probe the physics potential of IAXO in reconstructing quantities relevant for SN axion emission, if a detection is achieved.
Figure 7: $\chi^{2}$ as a function the different likelihood analysis parameters: $g_{a\gamma}^{2}g_{aN}^{2}$ (left) and the fraction of pions $\delta$ (right).
Figure 8: Likelihood analysis results in the $g_{a\gamma}^{2}g_{aN}^{2}$ and $\delta$ parameter space, in which the color palette shows the value of the $\chi^{2}$ distribution. The red dot represents the truth value of the SN axion model used as input for the calculations, while the white marker shows the minimum of the likelihood analysis in which the error bars correspond to one standard deviation ($\sigma$).

### v2

@@ 1 Introduction
Given the current technology capabilities, the detection of axions from a nearby SN event is a realistic possibility and the construction of a dedicated MeV detector for the next generation of axion helioscopes is highly motivated. First, a sensitivity to SN axions would allow us to explore regions of the axion parameter space not accessible through solar observations. Second, such a detection could shed light on important aspects of the SN physics, not accessible through other experimental tests. SNe are very rare events with current estimates predicting a rate of 1.63±0.461.63$\pm$ 0.461.63 ± 0.46 per century [14]. It is, therefore, imperative to ensure the highest level of preparedness for such an eventuality.
@@ 2 Supernova axion production
If axions are coupled to nuclear matter, the main production channels in the hot and dense SN core depend on their coupling to nucleons $g_{aN}$, with $SNaxionN=n,\,p$ for neutrons and protons respectively [18, 19, 20]. The first considered channel for axion emission from the SN nuclear medium is $NN$ bremsstrahlung $N+N\rightarrow N+N+a$ [21, 22, 23, 24, 25]. As depicted in the left panel of Fig. 1, in this process an axion is produced as a consequence of strong interactions between two nucleons in the medium. The state-of-the-art calculation for the bremsstrahlung emission rate has been performed in Ref. [26] and accounts for corrections beyond the usual one-pion exchange (OPE) approximation for the nuclear interaction potential. In particular, a non-vanishing mass for the exchanged pion [27], the contribution from the two-pion exchange [28], effective in-medium nucleon masses and multiple nucleon scattering effects [29, 30] have to be taken into account to obtain a reliable estimation of the bremsstrahlung emissivity. Finally, Ref. [31] has discussed the impact of finite-density effects over ALP-nucleon couplings in dense environments. Furthermore, the right panel of Fig. 1 displays a schematic representation of pionic Compton-like processes $\pi+N\rightarrow a+N$, in which a pion is converted into an axion after the scattering onto a non-relativistic nucleon. Since in the SN nuclear medium the fraction of thermal pions $Y_{\pi}$ was expected to be extremely small $Y_{\pi}\sim\mathcal{O}(10^{-4})$ [32], the contribution from pion conversions has been overlooked for a long time. Nevertheless, the authors of Ref. [33] pointed out that strong interactions can magnify the abundance of negatively-charged pions. Thus, the pion conversion emission rate has been re-estimated in Ref. [34], realizing that, under the assumption of Ref. [33], it could be comparable and even dominant over the bremsstrahlung contribution. Moreover, it has been recently shown that the contribution from the contact interaction term [35] and the $\Delta(1232)$ resonance [36] can significantly enhance axion emissivity via pion conversion. We address the reader to Refs. [18, 37] for the complete expressions of both $NN$ bremsstrahlung and pion conversion emission rates.
@@
If axions are weakly coupled to nuclear matter, ${g_{aN}\lesssim 10^{-8}}$, they are in the free-streaming regime, in which they can leave the star without being reabsorbed by the nuclear medium in the SN core. Therefore, they can act as an additional energy-loss channel during the SN cooling phase, contributing to the shortening of the neutrino burst duration [39]. In particular, consistency with observations of the neutrino burst from SN 1987A [^1] Current state-of-the-art SN simulations, including convection and updated neutrino-nucleon opacities, are in tension with late-time SN 1987A events, since they predict a cooling time shorter than the observed one [40]. requires ${g_{ap}\lesssim 5\times 10^{-10}}$ [19]. On the other hand, for sufficiently large couplings $g_{aN}\gtrsim 10^{-8}$, the SN core becomes “optically thick” for the produced axions, so that they can be also reabsorbed inside the core by means of the inverse processes $N+N+a\rightarrow N+N$ and $a+N\rightarrow\pi+N$ [19]. In this regime, dubbed trapping regime, axions are not able to escape the inner regions of the core and their emission can be approximately described as a blackbody radiation from a last-scattering surface [41], in close analogy with the case of neutrinos. We highlight that a complete treatment of the trapping regime would require the self-consistent addition of axions in SN simulations, since they could significantly impact the SN evolution. Relying just on the SN cooling argument, the range $5\times 10^{-10}\lesssim g_{ap}\lesssim 3\times 10^{-6}$ is excluded. However, larger values of the coupling can be constrained by the non observation of axion-induced events in Kamiokande-II during SN 1987A [42, 19]. Nevertheless, despite these constraints, even the trapped region remains a subject of interest for axion experiments [43, 44, 45, 46, 19, 17].
@@
In the following, we adopt as benchmark SN model the 1D spherical-symmetric GARCHING group’s SN model SFHo-s18.8 provided in Ref. [38], used also in previous studies about SN axions (see, e.g. Refs. [47, 48, 49, 18, 19, 50, 20]). The simulation, based on the neutrino-hydrodynamics code PROMETHEUS-VERTEX [51], is launched from a stellar progenitor with mass $18.8~M_{\odot}$ [52] and leads to a neutron star (NS) with baryonic mass $1.35~M_{\odot}$, a typical value expected for a SN explosion event [15]. More details on the employed SN model are provided in Appendix A. The presence of pions in the SN core has been taken into account by following the procedure illustrated in Ref. [53], including the pion-nucleon interaction as described in Ref. [33]. We highlight that the SN model employed in this work is characterized by relatively low SN core temperatures compared to the other profiles provided in Ref. [38] (see, e.g., Ref. [48, 54]). Therefore, it is expected to provide the most conservative results for the SN axion fluxes. A discussion of typical uncertainties related to SN limits on axion couplings is reported in Ref. [19]. Figure 2 shows the axion emission spectrum integrated over $10\,\operatorname{s}$ in both the free-streaming (black lines) and trapping regimes (red line) for our benchmark SN model. In the free-streaming regime, the spectrum is characterized by a bimodal shape, since the emission peak associated with the $NN$ bremsstrahlung is located at axion energies $E_{a}\sim 50\,\operatorname{MeV}$, while the pion conversion contribution peaks at $E_{a}\sim 200\,\operatorname{MeV}$ [18, 19]. This is related to the different nature of the two processes. In particular, the bremsstrahlung behaves as a quasi-thermal process, so that the average energy of the emitted ALP can be estimated as $E_{a}\simeq\frac{3}{2}T$  [50]. On the other hand, the pion conversion is highly non-thermal and the average ALP energy is given by $E_{a}\simeq\,m_{\pi}+\frac{3}{2}T$, where $m_{\pi}\simeq 135\,\operatorname{MeV}$ is the pion mass.
@@ 3 Direct detection of SN axions in IAXO
7 HD 157244 $\beta$ Arae 0.219±0.0100.219$\pm$ 0.0100.219 ± 0.010 17:25:17.99 -55:31:47.57
[…]
@@
It is evident that the detection of SN axions using IAXO and BabyIAXO relies heavily on our ability to predict core-collapse SN events with sufficient advance notice and angular precision to plan observations effectively. For this purpose, one has to rely on an early-alert system. Effectively, this is possible through the observation of pre-SN neutrinos, as the neutrino production rises rapidly during the last stages of nuclear burning [58]. Although the flux remains substantially lower—and at lower energies—than what is expected during the explosion itself, these neutrino fluxes are predicted to be observable in dedicated detectors minutes to days before the rise of the electromagnetic emission, providing the desired early alert (see, e.g., Ref. [58]). Dedicated analyses indicate, for example, that the combined analysis of KamLAND and Super-Kamiokande (SK) can detect the pre-SN neutrino flux from stars up to $\sim$ 500 pc, providing an early warning up to 12$\sim$ 12∼ 12 hours before the SN event [59]. Extracting directional information is more challenging. A detailed study of the potential to accurately detect the position of the source from pre-SN observations was presented in Ref. [55]. Information about the upcoming event, including positional data, is expected to be available hours before the event, a time more than sufficient to turn the IAXO telescope towards the star. The directional sensitivity of these events is relatively low, due to the weak forward-backward asymmetry of the inverse beta decay process used to detect the neutrino events. For current liquid scintillators, the low value of this asymmetry translates into a moderate angular resolution of around $60^{\circ}$ when 200 events are detected, which is realistic for a star like Betelgeuse. The use of a lithium-loaded liquid scintillator (LS-Li) would substantially increase the forward-backward asymmetry, pushing the angular resolution to about $15^{\circ}$ with the same number of events. [^3] Although currently there is no existing or funded large-volume LS-Li detector specifically optimized for pre-SN neutrino detection, ongoing developments indicate their feasibility. Experiments like PROSPECT [60] have successfully demonstrated lithium-loaded scintillator technology at multi-ton scales. Additionally, LS-Li detectors are actively being explored in reactor monitoring and neutrino oscillation experiments, such as the Segmented Anti-Neutrino Directional Detector (SANDD) project [61]. Furthermore, the conceptual design of the THEIA [62] detector includes lithium loading among its potential future upgrades, although it remains in a preliminary proposal stage. Finally, upcoming large-scale neutrino detectors, including DUNE Phase II [63] and potentially THEIA, are expected to significantly enhance angular resolution capabilities within the operational timeline of IAXO. Further improvements are possible with proposed future neutrino detectors, such as THEIA [62].
@@
Despite the moderate angular precision expected in the case of pre-SN neutrino events, the information provided by an early warning system may suffice to identify the SN candidate, given that the sample of possible close-by candidates is known (see Tab. 1), and that their relative angular distances are in most cases relatively large (cf. Tab. 5 in Appendix B). If the progenitor were correctly identified, (Baby)IAXO would have sufficient pointing accuracy to observe it during the collapse. Indeed, the pointing accuracy required for IAXO to follow the Sun—its primary purpose—is $\leq 30^{\prime\prime}$ and the field of view (FOV) associated with a 60–70 cm diameter magnet bore and detector is more than sufficient to enable accurate pointing toward any of the supernova candidates. [^4] Note that the most stringent pointing requirements for solar axion searches are given by the X-ray telescopes of IAXO. However, such telescopes are not available for MeV photons. Therefore, assuming the detector covers the entire bore, the FOV in this case is estimated from the diameter of the bore divided from the distance of the detector from the bore opening. For standard IAXO geometries, this corresponds to 24$\approx$ 2-4≈ 2 - 4 degrees. A more critical issue would arise if multiple potential targets fell within the angular uncertainty cone. Given the typically large angular separations between supernova sources relative to the IAXO FOV, a decision would be required in such cases. Since pointing to the wrong object could result in a significant loss of sensitivity, the most natural choice would be to point toward the nearest target, which is expected to yield the strongest signal. However, a definitive decision-making protocol—adapted to case-specific circumstances—has not been formally discussed or approved by the IAXO collaboration yet.
@@ 4 Sensitivity prospects
$S_{\gamma}=g_{a\gamma}^{2}\,g_{aN}^{2}\int_{E_{i}}^{E_{f}}\mathrm{d}E\,\frac{\mathrm{d}n_{\gamma}}{\mathrm{d}E}\leq\operatorname{ln}(20)\,.$ (4.5)
Sensitivity prospects are shown in Fig. 5 and 6 for two different SN progenitor candidates, namely Spica ($\alpha$ Virginis) and Betelgeuse ($\alpha$ Orionis). [^6] We underline that the value of the ALP-nucleon coupling assumed in the trapping regime, $g_{aN}=3\times 10^{-6}$, is disfavored, as it would have led to axion-induced events in Kamiokande II coinciding with the SN 1987A neutrino burst (see Refs. [42, 19]). Nevertheless, searches for supernova axions with IAXO would still offer a valuable and complementary probe of this region of parameter space. These sensitivities have been computed in different helioscope scenarios: BabyIAXO, consisting of two magnetic bores, with a length $L=10$ m and an area $A=0.77$ m2 as well as an average magnetic field $B=2$ T; IAXO, which will be equipped with 8 magnet bores with $L=20$ m and $A=2.3$ m2, providing a magnetic field of strength $B=2.5$ T; and IAXO+, which represents an enhanced version of IAXO, with a magnetic field $B=3.5$ T, an area $A=3.9$ m2 and magnet bores with $L=22$ m. These helioscope scenarios are described in Ref. [9] and summarized in Table 4. For the high-energy gamma detector we assume a conservative detection efficiency of $\epsilon=95\%$ and an energy threshold of 3 MeV.
@@
Figure 5: Sensitivity prospects for BabyIAXO, IAXO and IAXO+ as SNScopes for Betelgeuse (violet, left panel) and Spica (green, right panel) in the ALP free streaming regime considering bremsstrahlung emission and both bremsstrahlung and pion conversion emission. In the upper panels, we set the ALP-nucleon coupling to ${g_{aN}=10^{-9}}$, while in the lower panels we set it to the maximum value allowed by the SN cooling argument ${g_{aN}=5\times 10^{-10}}$. The black transparent regions show the sensivity prospects of IAXO, BabyIAXO and IAXO+ as helioscopes.
@@
Instead of using the natural likelihood, it is more convenient to work with the logarithm of the likelihood $\ln{L}$. According to Wilks’ theorem [67], if certain general conditions are satisfied, the minimum of $-2\ln{L}$ asymptotically approaches a $\chi^{2}$ distribution
@@
The likelihood analysis is performed for a fixed axion mass in the coherence region ($m_{a}=0.01$ eV), using Eq. (4.8). Here, $n_{i}$ represents the number of data counts for a given energy bin and $s_{i}$ represents the number of the expected counts obtained by integrating Eq. (4.7) within the same energy bin, scanning over the $g_{a\gamma}^{2}\times g_{aN}^{2}$ and $\delta$ parameter space. Mock data are generated using random energy counts distributed according to the free streaming input distribution shown in Fig. 2 with $\delta=1$ and a coupling constant of $g_{aN}=5\times 10^{-10}$. Fig. 7 and Fig. 8 show an illustrative example of the likelihood analysis results for 10 detected counts. The minimum of the $-2\ln{L}$ distribution is extracted, in which the shift of one $-2\ln{L}$ unit corresponds to one standard deviation ($\sigma$) in the parameter estimation. As shown, the values of the best fit parameters are close to the true values. These results probe the physics potential of IAXO in reconstructing quantities relevant for SN axion emission, if a detection is achieved.
Figure 7: The $-2\ln{L}$ distribution as a function the different likelihood analysis parameters: $g_{a\gamma}^{2}g_{aN}^{2}$ (left) and the fraction of pions $\delta$ (right).
Figure 8: Likelihood analysis results in the $g_{a\gamma}^{2}g_{aN}^{2}$ and $\delta$ parameter space, in which the color palette shows the value of the $-2\ln{L}$ distribution. The red dot represents the truth value of the SN axion model used as input for the calculations, while the white marker shows the minimum of the likelihood analysis in which the error bars correspond to one standard deviation ($\sigma$).